An orchestra has 30 musicians. Twelve of them can play the flute and twelve of them can play the trumpet. Six of them can play both. How many of the musicians can not play either of these instruments?
Answer by richwmiller(17219)

12-6=6
6+6+6=18
There are a total of 18 who play flute and or trumpet.
so there are 12 who play neither.
30-18=12